OpenStudy (mef4.0):

find y' y = (2x^7-x^2) ((x-1)\(x+1))

OpenStudy (anonymous):

ick product plus quotient rule use \[(fg)'=f'g+g'f\] with \[f(x)=2x^7-x^2,f'(x)=14x^6-2x, g(x) =\frac{x-1}{x+1},g'(x)=\frac{2}{(x+1)^2}\] put it all together to get your answer

OpenStudy (mef4.0):

Can you explain how you get the 2 in the numerator of g′(x)=2/(x+1)^2. I don't understand. And when you say put it all together, should I continue by multiplying it all out or just leave everything in parenthesis?